# Split Matching for Inductive Zero-shot Semantic Segmentation

### v2

@@ 3.2 Split Matching (SM)
To apply Hungarian matching in semantic segmentation, we need to split the ground-truth into class labels and their corresponding class-agnostic masks. We then compute the similarity between the predicted classes and masks and the ground-truth annotations. Based on these similarities, each ground-truth instance is assigned to a query and its associated mask. However, in ZSS, unannotated regions lack both ground-truth masks and class labels, making direct assignment infeasible. To address this issue, SM generates pseudo masks for latent classes (unseen candidates) in the unannotated areas and assigns them to candidate queries, while optimizing seen and candidate queries separately.
@@
To generate the pseudo masks and their corresponding class embeddings for unseen candidates, we use the multi-scale K-means and mask fusion methods from CLIP-ZSS [Chen et al.(2023)Chen, Deguchi, Zhang, Zheng, and Murase]. Specifically, given an image X, we feed it into a frozen CLIP visual encoder to obtain the CLIP dense features O. Then, we compute the seed for K-means methods with,
@@
where $I=\{0,[s/2],\dots,[H-s]\}$ and $J=\{0,[s/2],\dots,[W-s]\}$ are index sets, []delimited-[][$\cdot$][ ⋅ ] denotes rounding, and $s\in S$ is the window size. After we obtain G, we merge the masks corresponding to G by the mask fusion algorithm [Chen et al.(2023)Chen, Deguchi, Zhang, Zheng, and Murase], which merge the semantically similar masks by the cosine similarity among cluster centroids, to obtain pseudo label $\textbf{Y}_{u}\in[0,1]^{U\times H\times W}$ for unseen candidates where $U$ indicates the number of unseen candidates. After obtaining $\textbf{Y}_{u}$, we use $\textbf{Y}_{u}$ to mask the input images X into $\textbf{X}_{m}\in\mathbb{R}^{U\times 3\times H\times W}$. Finally, we feed $\textbf{X}_{m}$ into the frozen CLIP visual encoder for $\textbf{C}_{u}\in\mathbb{R}^{U\times C}$ as class embeddings for unseen candidates.
After obtaining $\textbf{Y}_{u}$ and $\textbf{C}_{u}$, we apply split matching as illustrated in Fig. 4. Specifically, we first feed the randomly initialized queries Q into a transformer decoder to obtain the predictions for each query, denoted as $\textbf{P}=\{\textbf{P}_{s},\textbf{P}_{u}\}$. Here, we denote the predictions for seen and candidate queries as $\textbf{P}_{s}=\{\textbf{V}_{s},\textbf{M}_{s}\}$ and $\textbf{P}_{u}=\{\textbf{V}_{u},\textbf{M}_{u}\}$, respectively. Specifically, $\textbf{V}_{s}\in\mathbb{R}^{K_{s}\times C}$ and $\textbf{M}_{s}\in\mathbb{R}^{K_{s}\times H\times W}$ denote the features after transformer decoder and predicted masks for the $K_{s}$ seen queries. Similarly, $\textbf{V}_{u}\in\mathbb{R}^{K_{u}\times C}$ and $\textbf{M}_{u}\in\mathbb{R}^{K_{u}\times H\times W}$ correspond to the predictions for the $K_{u}$ candidate queries. In both cases, the queries are projected into a $C$-dimensional semantic space. Similar to the conventional hungarian matching, SM also consists of class matching and mask matching. For class matching, we concatenate the seen class embeddings $\textbf{A}_{s}$ (extracted from the CLIP textual encoder) and candidate-class embeddings $\textbf{C}_{u}$ to form a joint class embedding $\textbf{E}=\mathrm{cat}(\textbf{A}_{s},\textbf{C}_{u})\in\mathbb{R}^{(N_{s}+U)\times C}$, where ‘$\mathrm{cat}$’ denotes concatenation. Finally, we separately compute the similarity between the predictions and joint class embeddings for seen and candidate queries, respectively. $\textbf{S}_{s}=Sigmoid(\textbf{V}_{s}\cdot\textbf{E}^{\top}),\textbf{S}_{u}=Sigmoid(\textbf{V}_{u}\cdot\textbf{E}^{\top})$. Although both seen and candidate queries compute similarity against the unified embedding matrix $\mathbf{E}=\text{cat}(\mathbf{A}_{s},\mathbf{C}_{u})$, we apply Hungarian matching separately for seen and candidate queries. This allows each group to focus on relevant candidates while still benefiting from global semantic context.
@@
$\sigma_{s}^{*}=\arg\min\sum_{i=0}^{K_{s}-1}\mathcal{L}_{\text{match}}(\textbf{S}_{s}^{\sigma(i)},\textbf{M}_{s}^{\sigma(i)},\textbf{E}_{i},\textbf{Y}_{s}^{i}),\quad\sigma_{u}^{*}=\arg\min\sum_{i=0}^{K_{u}-1}\mathcal{L}_{\text{match}}(\textbf{S}_{u}^{\sigma(i)},\textbf{M}_{u}^{\sigma(i)},\textbf{E}_{i},\textbf{Y}_{u}^{i})$ (2)
@@
After computing $\sigma_{s}^{*}$ and $\sigma_{u}^{*}$, we concatenate them into a unified assignment $\sigma^{*}$. The resulting assignment $\sigma^{*}$ is then used to optimize the matching loss across all queries. $\mathcal{L}_{match}$ is composed of two terms: a classification loss $\mathcal{L}_{cls}$ and a mask loss $\mathcal{L}_{mask}$:
$\mathcal{L}_{\text{match}}(\textbf{S},\textbf{M},\textbf{E},\textbf{Y})=\mathcal{L}_{cls}(\textbf{S}^{\sigma^{*}},\textbf{E})+\mathcal{L}_{mask}(\textbf{I}^{\sigma^{*}},\textbf{Y}),$ (3)
@@
where $\textbf{S}\in\mathbb{R}^{K\times C}$ denotes classification scores over $C$ classes for $K$ queries, and $\textbf{M}\in\mathbb{R}^{K\times H\times W}$ represents the corresponding predicted masks. $\textbf{S}^{\sigma^{*}}$ and $\textbf{I}^{\sigma^{*}}$ are the classification scores and mask similarities after matching with the optimal assignment $\sigma^{*}$. To enhance the discriminability of candidate queries, we introduce a cosine similarity loss: $\mathcal{L}_{cos}=1-\mathrm{cos}(\textbf{V}_{u}^{\prime},\textbf{C}_{u})$, where $\textbf{V}_{u}^{\prime}$ are the matched candidate queries, and $\mathrm{cos}(a,b)$ denotes cosine similarity. The final training objective for split matching is $\mathcal{L}_{SM}=\mathcal{L}_{match}+\mathcal{L}_{cos}$.
3.3 Multi-scale Feature Enhancement
@@
While SM effectively enables query-based approaches to tackle ZSS, the associated keys and values in transformer decoders used to locate corresponding regions in an image remain suboptimal, yet the current design lacks sufficient refinement, limiting the performance of the matching. To tackle this issue, we propose an Multi-scale Feature Enhancement (MFE) module, as illustrated in Fig. 4, designed to assist in identifying the relevant regions by effectively combining multi-scale features. The MFE leverages hierarchical features extracted by the pixel decoder to provide a comprehensive representation, capturing both fine-grained and global contextual information. The multi-scale outputs of the pixel decoder, $\textbf{F}=\left\{\textbf{F}_{0},\textbf{F}_{1},\textbf{F}_{2}\right\}$, represent features at finer resolutions, where $\textbf{F}_{0}$ has the coarsest resolution and $\textbf{F}_{2}$ the finest where $\textbf{F}_{i}\in\mathbb{R}^{C\times(H/r^{2-i})\times(W/r^{2-i})}$, with $r$ denoting the scale factor.
@@
The fusion process begins by refining $\textbf{F}_{0}$, the lowest-resolution feature map, through a dense block consisting of a convolutional layer, group normalization [Wu and He(2018)], and a ReLU activation function. This block enhances local spatial relationships and prepares the feature map for integration with features at finer scales. The refined $\textbf{F}_{0}$ is then resized to align with the spatial dimensions of $\textbf{F}_{1}$, which undergoes a similar refinement process. The refined $\textbf{F}_{0}$ and $\textbf{F}_{1}$ are fused via element-wise addition, combining detailed lowest-resolution features with mid-level semantic information. This combined representation is resized again to match the spatial dimensions of $\textbf{F}_{2}$, the finest feature map in the hierarchy. Simultaneously, $\textbf{F}_{2}$ is processed through its own dense block to extract refined features. The previously fused representation of $\textbf{F}_{0}$ and $\textbf{F}_{1}$ is then combined with the refined $\textbf{F}_{2}$ to produce the final unified feature map, $\textbf{F}_{u}$. This hierarchical fusion ensures that fine-grained details from the highest-resolution features are effectively integrated with broader contextual cues. The final output, $\textbf{F}_{u}$, leverages complementary information across scales to provide a rich, unified representation that allows queries to accurately identify relevant regions in the image. After obtaining $\textbf{F}_{u}$, we optimize it through $\mathcal{L}_{MFE}=\mathcal{L}_{ce}(\textbf{F}_{u},\textbf{Y}_{p})+\mathcal{L}_{focal}(\textbf{F}_{u},\textbf{Y}_{p})$ where $\mathcal{L}_{ce}$ and $\mathcal{L}_{focal}$ are cross entropy and focal loss [Lin et al.(2017)Lin, Goyal, Girshick, He, and Dollár].
@@ 3.4 Training and Inference
The total loss functions are $\mathcal{L}=\mathcal{L}_{SM}+\mathcal{L}_{MFE}$. During inference, we exclude the MFE module and utilize only the backbone and transformer decoder. Both trained seen and candidate queries, along with additional Randomly-initialized Queries (RQ) $\textbf{Q}_{r}$, are used to produce the final prediction. These RQs increase the density of queries in the feature space, enabling better coverage of unannotated regions and discovering additional latent candidates. The inference process is further illustrated in the Supplementary Materials.
@@ 4 Experiments
Dataset. To assess the effectiveness of our proposed method, we conduct experiments on the widely-used benchmark COCO-Stuff [Caesar et al.(2018)Caesar, Uijlings, and Ferrari] and PASCAL VOC [Everingham et al.(2015)Everingham, Eslami, Van Gool, Williams, Winn, and Zisserman], focusing on the task of zero-shot semantic segmentation (ZSS). We adopt the same seen and unseen category splits as in prior works [Ding et al.(2022)Ding, Xue, Xia, and Dai, Zhou et al.(2023)Zhou, Lei, Zhang, Liu, and Liu, Zhou et al.(2022)Zhou, Loy, and Dai]. Specifically, COCO-Stuff consists of a total of 171 classes with 156 seen and 15 unseen classes according to the standard protocol. The dataset includes 118,287 images for training and 5,000 images for testing. PASCAL VOC contains 10,582 images for training and 1,449 images for validation, including 15 seen and 5 unseen classes.
      Models       Backbone       PASCAL VOC       COCO-Stuff
@@
      SPNet [Xian et al.(2019)Xian, Choudhury, He, Schiele, and Akata]       ResNet101 [He et al.(2016)He, Zhang, Ren, and Sun]       26.1       78.0       15.6       14.0       35.2       8.7
@@
      Zzseg [Xu et al.(2022)Xu, Zhang, Wei, Lin, Cao, Hu, and Bai]       ViT-B [Dosovitskiy et al.(2020)Dosovitskiy, Beyer, Kolesnikov, Weissenborn, Zhai, Unterthiner, Dehghani, Minderer, Heigold, Gelly, et al.]       77.5       83.5       72.5       37.8       39.3       36.3
@@
      ZegCLIP [Zhou et al.(2023)Zhou, Lei, Zhang, Liu, and Liu]       84.3       91.9       77.8       40.8       40.2       41.4
      OTSeg [Ye et al.(2024)Ye, Oh, et al.]       84.5       92.1       78.1       41.4       41.4       41.4
      Ours       ResNet101 [He et al.(2016)He, Zhang, Ren, and Sun]       85.3       87.7       83.1       42.5       42.6       42.4
@@
Implementation Details. The CLIP model applied in our method is based on the ViT-B/16 model, and the channel of the output text features is 512. All the experiments are conducted on 8 V100 GPUs, and the batch size is set to 16 for both datasets. The iterations are set to 20K and 80K for PASCAL VOC and COCO-Stuff. $\mathcal{L}_{cls}$ in $\mathcal{L}_{match}$ is focal loss [Lin et al.(2017)Lin, Goyal, Girshick, He, and Dollár] and the $\mathcal{L}_{mask}$ is a combination of IoU loss and DICE loss [Cheng et al.(2022)Cheng, Misra, Schwing, Kirillov, and Girdhar] in $\mathcal{L}_{match}$. We choose Mask2Former [Cheng et al.(2022)Cheng, Misra, Schwing, Kirillov, and Girdhar] with ResNet101 as the backbone with all other hyperparameters unchanged. 50 unseen and 50 random queries are added during inference. We apply the harmonic mean IoU (hIoU) following previous works [Zhou et al.(2023)Zhou, Lei, Zhang, Liu, and Liu] where $hIoU=\frac{2\cdot sIoU\cdot uIoU}{sIoU+uIoU}.$ as the metric where $sIoU$ and $uIoU$ indicate the mIoU of the seen classes and unseen classes, respectively. More details are in the Supplementary Materials.

### v3

@@ 3.2 Split Matching (SM)
Hungarian matching requires splitting ground-truth into class labels and corresponding class-agnostic masks, followed by assigning each ground-truth instance to a query based on similarity. However, in ZSS, unannotated regions lack ground-truth labels, making it impossible to directly apply Hungarian matching, which relies on full supervision. To overcome this, Split Matching (SM) generates pseudo masks for latent classes in unannotated areas and assigns them to candidate queries, while optimizing seen and candidate queries separately.
@@
To generate the pseudo masks and their corresponding class embeddings for unseen candidates, we use the multi-scale K-means and mask fusion methods from CLIP-ZSS [Chen et al.(2023)Chen, Deguchi, Zhang, Zheng, and Murase]. Specifically, given an image X, we feed it into a frozen CLIP visual encoder to obtain the CLIP dense features O. Then, we compute the seed for K-means methods with
@@
where $I=\{0,[s/2],\dots,[H-s]\}$ and $J=\{0,[s/2],\dots,[W-s]\}$ are index sets, $\cdot$] denotes rounding, and $s\in S$ is the window size. After we obtain G, we merge the masks corresponding to G by the mask fusion algorithm [Chen et al.(2023)Chen, Deguchi, Zhang, Zheng, and Murase], which merge the semantically similar masks by the cosine similarity among cluster centroids, to obtain pseudo label $\textbf{Y}_{u}\in[0,1]^{U\times H\times W}$ for unseen candidates where $U$ indicates the number of unseen candidates. After obtaining $\textbf{Y}_{u}$, we use $\textbf{Y}_{u}$ to mask the input images X into $\textbf{X}_{m}\in\mathbb{R}^{U\times 3\times H\times W}$. Finally, we feed $\textbf{X}_{m}$ into the frozen CLIP visual encoder for $\textbf{C}_{u}\in\mathbb{R}^{U\times C}$ as class embeddings for unseen candidates. Unlike [Amir et al.(2021)Amir, Gandelsman, Bagon, and Dekel] which rely on DINO-ViT [Caron et al.(2021)Caron, Touvron, Misra, Jégou, Mairal, Bojanowski, and Joulin] features, saliency voting, and post-processing (GrabCut), our method directly forms candidate masks without voting or additional refinement. Compared with [Rewatbowornwong et al.(2023)Rewatbowornwong, Chatthee, Chuangsuwanich, and Suwajanakorn], we do not fix the number of clusters or use image-to-text captioning with CLIP, ensuring no leakage of unseen semantics and reducing computational overhead.
@@
$\sigma_{s}^{*}=\arg\min\sum_{i=0}^{K_{s}-1}\mathcal{L}_{match}(\textbf{S}_{s}^{\sigma(i)},\textbf{M}_{s}^{\sigma(i)},\textbf{E}_{i},\textbf{Y}_{s}^{i}),\quad\sigma_{u}^{*}=\arg\min\sum_{i=0}^{K_{u}-1}\mathcal{L}_{match}(\textbf{S}_{u}^{\sigma(i)},\textbf{M}_{u}^{\sigma(i)},\textbf{E}_{i},\textbf{Y}_{u}^{i})$ (2)
@@
where $\mathcal{L}_{match}(\textbf{S}^{\sigma},\textbf{M}^{\sigma},\textbf{E},\textbf{Y})=\mathcal{L}_{cls}(\textbf{S}^{\sigma},\textbf{E})+\mathcal{L}_{mask}(\textbf{I}^{\sigma},\textbf{Y})$, consisting of $\mathcal{L}_{cls}$ and $\mathcal{L}_{mask}$. The classification loss $\mathcal{L}_{cls}$ implemented using focal loss, while the mask loss $\mathcal{L}_{mask}$ is computed using the DICE loss. After computing $\sigma_{s}^{*}$ and $\sigma_{u}^{*}$, we concatenate them into a unified assignment $\sigma^{*}$. $\sigma^{*}$ is then used to optimize the matching loss $\mathcal{L}^{*}_{match}$ across all queries,
$\mathcal{L}^{*}_{match}(\textbf{S},\textbf{M},\textbf{E},\textbf{Y})=\mathcal{L}_{cls}(\textbf{S}^{\sigma^{*}},\textbf{E})+\mathcal{L}_{mask}(\textbf{I}^{\sigma^{*}},\textbf{Y}_{p}),$ (3)
@@
where $\textbf{S}^{\sigma^{*}}$ and $\textbf{I}^{\sigma^{*}}$ are the classification scores and mask similarities after matching with the optimal assignment $\sigma^{*}$. To enhance the discriminability of candidate queries, we introduce a cosine similarity loss: $\mathcal{L}_{cos}=1-\mathrm{cos}(\textbf{V}_{u}^{\prime},\textbf{C}_{u})$, where $\textbf{V}_{u}^{\prime}$ is the candidate queries which are assigned with non-igored classes under $\sigma^{*}_{u}$, and $\mathrm{cos}(a,b)$ denotes cosine similarity. The final loss for split matching is $\mathcal{L}_{SM}=\mathcal{L}^{*}_{match}+\mathcal{L}_{cos}$. Importantly, $\mathcal{L}_{cos}$ is not merely an auxiliary term, but a necessary component of the Split Matching framework, ensuring stable and discriminative alignment of candidate queries in the absence of ground-truth.
3.3 Multi-scale Feature Enhancement (MFE)
@@
Although SM facilitates the adaptation of query-based approaches to zero-shot segmentation, the key and value features in transformer decoders, responsible for associating queries with relevant image regions, remain suboptimal. Due to the lack of further refinement, the matching performance suffers, ultimately constraining the full potential of the model. To tackle this issue, we propose an Multi-scale Feature Enhancement (MFE) module, as illustrated in Fig. 4, designed to assist in identifying the relevant regions by effectively combining multi-scale features. The MFE leverages hierarchical features extracted by the pixel decoder to provide a comprehensive representation, capturing both fine-grained and global contextual information. The multi-scale outputs of the pixel decoder, $\textbf{F}=\left\{\textbf{F}_{0},\textbf{F}_{1},\textbf{F}_{2}\right\}$, represent features at finer resolutions, where $\textbf{F}_{0}$ has the coarsest resolution and $\textbf{F}_{2}$ the finest where $\textbf{F}_{i}\in\mathbb{R}^{C\times(H/r^{2-i})\times(W/r^{2-i})}$, with $r$ denoting the scale factor.
@@
The fusion process begins by refining the lowest-resolution feature map $\textbf{F}_{0}$ using a dense block with a convolution, group normalization [Wu and He(2018)], and ReLU activation. This block strengthens local spatial cues and prepares $\textbf{F}_{0}$ for integration with finer-scale features. The refined $\textbf{F}_{0}$ is resized to match $\textbf{F}_{1}$, which is similarly refined, and the two are fused via element-wise addition to combine coarse and mid-level semantics. The fused representation is then resized to match $\textbf{F}_{2}$, the highest-resolution map, which is also refined through its dense block. Finally, the fused $\textbf{F}_{0}$ and $\textbf{F}_{1}$ are merged with refined $\textbf{F}_{2}$ to produce the unified feature map $\textbf{F}_{d}$, which integrates fine-grained details and contextual cues across scales. We optimize $\textbf{F}_{d}$ with $\mathcal{L}_{MFE}=\mathcal{L}_{ce}(\textbf{F}_{d},\textbf{Y}_{p})+\mathcal{L}_{focal}(\textbf{F}_{d},\textbf{Y}_{p})+\mathcal{L}_{global}(\textbf{F}_{d},\textbf{C}_{g})$, where $\mathcal{L}_{ce}$, $\mathcal{L}_{focal}$, $\mathcal{L}_{global}$ denote cross entropy, focal loss [Lin et al.(2017)Lin, Goyal, Girshick, He, and Dollár] and global loss in [Chen et al.(2023)Chen, Deguchi, Zhang, Zheng, and Murase].
@@ 3.4 Training and Inference
The total loss is $\mathcal{L}=\mathcal{L}_{SM}+\mathcal{L}_{MFE}$. During inference, we exclude the MFE module and utilize only the backbone and transformer decoder and propose a Random Query (RQ) strategy. Specifically, we feed the trained seen and candidate queries, along with new randomly initialized queries $\mathbf{Q}_{r}$, into the Transformer decoder. These queries collectively interact with the visual features to generate segmentation masks, where the $\mathbf{Q}_{r}$ serve to enrich query diversity and improve coverage of unannotated regions. Unlike the trained queries, $\mathbf{Q}_{r}$ are not supervised during training and are introduced only at inference time to probe unannotated or ambiguous regions in the image. By increasing the density and diversity of queries in the feature space, $\mathbf{Q}_{r}$ enhance the model’s ability to explore underrepresented regions that might correspond to unseen or latent objects. Importantly, $\mathbf{Q}_{r}$ act as complementary probes that are not biased by learned semantic categories, enabling the model to capture alternative activation patterns and recover instances that trained queries might overlook. More detailed inference process and the role of $\mathbf{Q}_{r}$ are illustrated in the Supplementary Materials.
@@ 4 Experiments
Dataset. We conduct experiments on the widely-used benchmark COCO-Stuff [Caesar et al.(2018)Caesar, Uijlings, and Ferrari] and PASCAL VOC [Everingham et al.(2015)Everingham, Eslami, Van Gool, Williams, Winn, and Zisserman], focusing on the task of zero-shot semantic segmentation (ZSS). We adopt the same seen and unseen class splits as in prior works [Ding et al.(2022)Ding, Xue, Xia, and Dai, Zhou et al.(2023)Zhou, Lei, Zhang, Liu, and Liu, Zhou et al.(2022)Zhou, Loy, and Dai]. Specifically, COCO-Stuff consists of a total of 171 classes with 156 seen and 15 unseen classes according to the standard protocol. The dataset includes 118,287 images for training and 5,000 images for testing. PASCAL VOC contains 10,582 images for training and 1,449 images for validation, including 15 seen and 5 unseen classes.
   Models    PASCAL VOC    COCO-Stuff
@@
   SPNet [Xian et al.(2019)Xian, Choudhury, He, Schiele, and Akata]    26.1    78.0    15.6    14.0    35.2    8.7
@@
   Zzseg [Xu et al.(2022)Xu, Zhang, Wei, Lin, Cao, Hu, and Bai]    77.5    83.5    72.5    37.8    39.3    36.3
@@
[…]
   Ours    85.3    87.7    83.1    42.5    42.6    42.4
@@
Implementation Details. The CLIP model applied in our method is based on the ViT-B/16 model, and the channel of the output text features is 512. All the experiments are conducted on 8 V100 GPUs, and the batch size is set to 16 for both datasets. The iterations are set to 20K and 80K for PASCAL VOC and COCO-Stuff. $\mathcal{L}_{cls}$ in $\mathcal{L}_{match}$ is focal loss [Lin et al.(2017)Lin, Goyal, Girshick, He, and Dollár] and the $\mathcal{L}_{mask}$ is a combination of IoU loss and DICE loss [Cheng et al.(2022)Cheng, Misra, Schwing, Kirillov, and Girdhar] in $\mathcal{L}_{match}$. We choose Mask2Former [Cheng et al.(2022)Cheng, Misra, Schwing, Kirillov, and Girdhar] with ResNet101 as the backbone with all other hyperparameters unchanged. 50 unseen and 50 random queries are added during inference. We apply the harmonic mean IoU (hIoU) following previous works [Zhou et al.(2023)Zhou, Lei, Zhang, Liu, and Liu] where $\small{hIoU}=\frac{2\cdot sIoU\cdot uIoU}{sIoU+uIoU}$ as the metric where $sIoU$ and $uIoU$ indicate the mIoU (mean intersection over union) of the seen classes and unseen classes, respectively. More details are in the Supplementary Materials.